Is (x+y)2 equal to (x2+y2)? Explain why or why not using an example to support your answer
I seem to be stuck :[

Found 2 solutions by josgarithmetic, richwmiller:
Answer by josgarithmetic(39613)

Depends what you mean. If you mean the expressions exactly as you wrote them, then YES. Reason: DISTRIBUTIVE PROPERTY. If you wrote those as you intended, then (x+y) is a factor, and 2 is a factor. Which is on the left or the right is not important (for real or complex numbers).

If your 2 is an exponent, then NO.

Answer by richwmiller(17219)

(x+y)^2= x^2+2xy+y^2
if x=0 or y=0 then (x+y)^2 = x^2+y^2
But any other values for x or y then they are not equal
x^2+y^2=(x+y)^2